### v1

# EE3P3D: Event-based Estimation of Periodic Phenomena Frequency using 3D Correlation

Abstract
We present a novel method for measuring the frequency of periodic phenomena, e.g., rotation, flicker and vibration, by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution. The approach assumes that for a periodic phenomenon, a highly similar set of events is generated within a specific spatio-temporal window at a time difference corresponding to the phenomenon’s period. The sets of similar events are detected by 3D spatio-temporal correlation in the event stream space. The proposed method, EE3P3D, is evaluated on a dataset of $12$ sequences of periodic phenomena, i.e. flashing light and vibration, and periodic motion, e.g., rotation, ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM). EE3P3D significantly outperforms published methods on this dataset, achieving the mean relative error of $0.1\%$.
Keywords:
Event camera Frequency estimation Periodic phenomena Frequency of periodic phenomena.
1 Introduction
Accurate measurement of periodic phenomena is important in various scientific and industrial fields. Precise quantification of rotational speed, for instance, finds applications in (i) monitoring rotating machinery components for performance evaluation and quality control [4], (ii) ensuring flight stability and manoeuvrability in drones [15], (iii) analysing sports equipment like ball tracking [9, 7], and (iv) optimising energy production in wind turbines [16].
Traditional methods for measuring periodic phenomena often involve contact-based devices such as tachometers and rotary encoders. These approaches, while established, have inherent limitations: (i) physical contact with the target object can influence its movement and introduce measurement inaccuracies, (ii) contact methods may not be feasible in scenarios with delicate objects, confined spaces, or situations where interfering with the target significantly affects its movement. Laser tachometers offer a less invasive and highly accurate [^1] See https://meters.uni-trend.com/product/ut370-series/#Specifications. alternative but require reflective material (e.g., a sticker) on the target and precise aiming of the laser, as missing the reflective material pass-through results in an inaccurate measurement. Although contactless event-based methods exist, they are either handcrafted for a specific type of periodic phenomena [2] or require user supervision [13, 14].
Figure 1: EE3P3D: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a 3D correlation of the template with the event stream is computed, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows.
In this paper, we propose a non-contact method for measuring the properties of any periodic phenomena using an event camera. The input of our method is the event stream, and the output is a single scalar – the frequency of the periodic phenomena. Our approach assumes that similar sets of pixel activations of an event camera (events) will occur at periodic time intervals in any sequence capturing a periodic phenomenon. The periodic intervals correspond to the target frequency.
We evaluated our method on a dataset of twelve sequences of four types of periodic phenomena: (i) light flickering, (ii) object vibration, (iii) rotational speed, and (iv) periodic movement and achieved a mean relative error of $0.1\%$, which is 140$\approx$ 140$\times$ 140 × lower than the mean relative error achieved by the best available published method (see Tab. 4).
2 Related work
In this section, we discuss existing approaches and technologies, mainly in the domain of rotation speed measurement, as this is the most commonly examined area of application.
Commercially available solutions for rotation speed measurement include contact and contactless methods. Contact-based devices include traditional mechanical tachometers that connect directly to a rotating shaft. These can introduce potential inaccuracies due to added mass and friction. Electrostatic sensors offer a contactless approach, detecting changes in the electromagnetic field around a rotating object (equipped with a bearing). The rotational speed is the frequency of these changes [11]. Optical encoder tachometers use a photoelectric sensor to detect light passing through a rotating disc with patterned segments. The rotation speed is the frequency of light intensity changes [1]. Finally, laser tachometers utilise a laser beam that bounces off reflective markers on the rotating target. The rotation speed is the frequency of the light reflections detected by a sensor.
Several camera-based methods for rotational speed measurement have been proposed. Wang et al. [18] proposed a frame-by-frame 2D correlation analysis for measuring rotation speeds up to $700$ RPM with relative error of ±$\pm 1\%$ using a low-cost RGB camera and a marker on the rotor. In [17], multiple image similarity methods were used to analyse similarities between consecutive frames and Chirp-Z transform to calculate the rotation speed, extending the measuring range to $900$ RPM. Natali et al. [12] further extended the measuring range by using 2D correlation and short-time Fourier Transform for measuring rotation speeds up to $1500$ RPM, highlighting the camera frame rate as the main limitation for measuring faster rotation. While being non-contact, RGB camera-based methods are often limited by the frame rate of the camera. Additionally, markers required by [17, 18] limit applicability.
The next group of methods utilise event cameras with a temporal resolution two orders of magnitude higher than RGB cameras. The EB-ASM method [2] measures rotation speeds up to $8500$ RPM by calculating the elapsed time between spikes of distinct polarity events caused by rotating patterns in a manually selected region. However, the method struggles when noise is present, and targets lack high-contrast features. In EV-Tach [19] method, only rotating objects of centrosymmetric shapes (e.g., propeller blades) are assumed, and the method allows for rotation speed measurements of up to $6000$ RPM with a relative mean absolute error as low as $0.3$‰. The Frequency-cam [13] method uses a second-order digital infinite impulse response (IIR) filter to reconstruct per-pixel brightness levels. The time deltas between zero-level crossing are then used to estimate the rotation speed.
In [7], a method for estimation of a table tennis ball spin was proposed. The method uses ordinal time surfaces for tracking, the spin of the ball is estimated from optical flow of events generated by a logo on the ball. The method achieved a spin magnitude mean error of $10.7\pm 17.3$ RPS and a spin axis mean error of $32.9\pm 38.2\degree$ in real time for a flying ball. In [10], accurate frequency estimation of periodic phenomena lacking high-contrast features was demonstrated using a 2D correlation of spatial patterns. However, their method requires manual selection of correlation areas, whereas in our proposed method, no user intervention is necessary.
We propose two additional baseline methods for evaluating the performance of our proposed method. Those two methods are inspired by the concepts presented in the works mentioned above and are described in the following section. Note that Prophesee, the manufacturer of event cameras, provides a method for estimating the frequency of periodic phenomena [14]. Although their approach is unknown, we present the results of their method in our experiments.
3 Methods
In this section, we describe three proposed methods. The Simple (Sec. 3.1) and FFT (Sec. 3.2) methods serve as baselines, and the proposed method is presented in Sec. 3.3.
3.1 A Simple baseline
A baseline method was implemented to test the hypothesis that separate per-pixel events analysis is insufficient for accurate frequency estimation of complex periodic phenomena. The event camera produces a sequence of tuples $(x,y,p,t)$, where $x,y$ are spatial coordinates, $p$ is event polarity (positive – $1$ or negative – 00), and $t$ is a timestamp in microseconds of the event creation.
For each pixel at position $x,y$, we construct a list of timestamps $T_{x,y,p}$ of a given polarity $p$.
The baseline method then proceeds in three steps:
  1. 1.
    Temporal analysis – the delta times between consecutive timestamps in $T_{x,y,p}$ are calculated, and their median is computed (temporal analysis result). This is performed separately for each pixel and polarity, producing $2\times W^{2}$ results per window size $W$.
  2. 2.
    Spatial analysis – temporal analysis results are aggregated using the median, which produces a single spatial analysis result per window, representing the estimated period of the phenomenon in microseconds.
  3. 3.
    Frequency prediction – the final predicted frequency for each window is derived from the estimated period $T$ in microseconds using the following equation $\nu(T)={10^{6}}/{T}$, where $\nu(T)$ returns frequency in Hertz (Hz).
3.2 A FFT baseline
The FFT baseline method estimates per-pixel frequency in four steps:
  1. 1.
    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform with the efficient Fast Fourier Transform (FFT) algorithm [3] with the timestamp of event creation represented as a Dirac pulse.
  2. 2.
    Signal smoothing – the Hanning window is applied before the Fourier transform.
  3. 3.
    Median filtering – a 2D median filter using a local window size of $3$ is applied to the 2D array of estimated per-pixel frequencies to mitigate outliers.
  4. 4.
    Results aggregation – the mode of detected frequencies is returned.
3.3 EE3P3D - The proposed method
Broadly speaking, the event stream is quantized and 3D correlation responses analysed to detect periodic similarities along the time axis of the event data. An overview of the method is shown in Fig. 1. We identify the period of a spatio-temporal pattern in the event stream in four steps:
  1. 1.
    Event quantization – a 3D spatio-temporal array $\mathbf{E}$ is created by aggregating sparse events within non-overlapping time intervals of length $t_{\text{quant}}$ into a zero-initialized array along the third axis. For each quantised interval, the polarity ($1$ for positive, $-1$ for negative) of the last event in that interval is recorded for each pixel.
  2. 2.
    Correlation template selection – the array $\mathbf{E}$ is split into a grid of $K$ spatially non-overlapping areas $\mathbf{A}_{i}$, $i\in\{1,\dots K\}$, of size $W\times W$. In each area, the depth (time axis size) of the correlation template of the area $\mathbf{T}_{i}$ is found by thresholding on a predefined number of events $N$.
  3. 3.
    3D correlation – a 3D correlation of the template $\mathbf{T}_{i}$ is computed with the area $\mathbf{A}_{i}$ (see Fig. 2). The time difference between consecutive peaks in correlation responses is the estimated period of the phenomenon.
  4. 4.
    Results aggregation – the output of the method is the median of estimated periods from all areas $\mathbf{A}_{i}$.
Figure 2: Normalised correlation responses of a selected 3D template with $1000$ ms of spatio-temporal event stream. Periodic peaks are highly distinctive, highly regular and indicate a periodic phenomenon.
4 Experiments
In this section, we discuss the event camera used and the selection of parameters for methods presented in the previous section, followed by experiments measuring rotation speed, frequency of periodic light flashes, vibrations, and moving objects.
In our experiments, we used the Prophesee EVK4 HD event camera with $1280\times 720$ resolution, capturing up to $1066$ million events per second [6] under light conditions ranging from $0.08$ up to 100 000100000100\,000100 000 lux [^2] Spec sheet available at https://bit.ly/Sony-Prophesee-IMX636-IMX637-specsheet. with microsecond time resolution (equivalent to 10 0001000010\,00010 000 frames per second).
The IMX636 sensor within the event camera allows fine-tuning its behaviour through five adjustable biases [^3] Described in detail at https://docs.prophesee.ai/stable/hw/manuals/biases.html.. Contrast sensitivity thresholds, controlled by bias_diff_on and bias_diff_off, determine how much brighter or darker the scene needs to become before a pixel generates an event. In our experiments, we set them to $50$ and $30$, respectively, due to a higher noise of events with positive polarities. Bandwidth biases, bias_fo and bias_hpf, control low-pass and high-pass filters, respectively, influencing how the sensor responds to rapid or slow light changes. Finally, the dead time bias (bias_refr) regulates the time of non-responsiveness of a pixel after generating an event. We left the remaining biases at their default settings.
We evaluated the accuracy of the Simple baseline method (see Sec. 3.1) when using window sizes $W$ of $30\times 30$ px, $45\times 45$ px, $60\times 60$ px and $75\times 75$ px and achieved the lowest mean relative error (MRE) by setting $W$ to $45\times 45$ px. Further increasing or decreasing window size or using mode instead of the median in the temporal or spatial analysis steps of this method negatively impacted the accuracy.
The FFT baseline method (see Sec. 3.2) achieved lower mean relative error when using timestamps of events with negative polarity (MRE of $60\%$ as opposed to $67.467\%$).
We evaluated all different configurations of parameters (window sizes $W$ of $30$, $45$, $60$ and $75$ and counts of template events $N$ of $100$, $200$, …, $3000$) for the proposed method (see Sec. 3.3). In all our experiments, the length of the quantisation window $t_{\text{quant}}$ is set to $100$ microseconds. When using window size $W$ set to $45\times 45$ px the method achieves the lowest MRE by a substantial margin and a low maximal relative error compared to results when using other window sizes (see Tab. 1). We noticed a significant improvement in accuracy when using counts of template events $N$ higher than $1500$ (see Tab. 2). The best accuracy was achieved using the window size $W$ set to $45\times 45$ px and the template event count $N$ set to $1800$ events. These parameters are proposed as default parameters of the method.
To objectively compare the results of the proposed method, we use the mentioned baseline methods (see Sec. 3.1, 3.2), a vibration estimation method [14] developed by Prophesee, manufacturer of the used event camera, the open-source Frequency-cam method [13] and re-implemented EB-ASM method [2]. As the output of Frequency-cam is in the form of colourmaps, we modified it to instead output a reading for each pixel that would be used for colourmap creation and computed their median. Since every method produces a different amount of measurements per one second of input data, we calculate the median of measurements to obtain a single value for a comparison with other methods.
Window size (px) $30\times 30$ $45\times 45$ $60\times 60$ $75\times 75$
Avg rel error $37.78$ $\mathbf{19.34}$ $42.96$ $487.98$
Max rel error $\mathbf{3225.00}$ $3362.50$ $7712.50$ $22737.50$
Table 1: EE3P3D - dependence on window size. Average and maximum relative errors (%) on the 12 sequence dataset.
Template event count $100-1400$ $1500-3000$
Average relative error $308.46$ $\mathbf{5.75}$
Maximum relative error $22737.50$ $\mathbf{269.36}$
Table 2: EE3P3D - dependence on template event count (see Sec. 3.3). Average and maximum relative errors (%) on the $12$ sequence dataset.
4.1 Dataset
To capture ground truth (GT) rotation speed data, the Uni-Trend UT372 laser tachometer [^4] Details at https://meters.uni-trend.com/product/ut370-series. was used. The tachometer range is $10$ to 99 $\,9999$ RPM with a relative error of ±$\pm 0.04\%$. We averaged all tachometer readings captured in one second ($3$ to $5$) to obtain a single reference value for comparison with the results of all methods.
Figure 3: Ground-truth frequencies of experiments in order of appearance in this work.
However, acquiring GT data this way was not feasible for all experiments. In such cases, the EE3P method [10] was used to analyse the data multiple times with varying parameter configurations to filter potential measurement outliers and obtain a GT frequency estimate as close as possible to the actual value. The frequency was then verified by manually examining the event data stream.
Event data for three experiments in this work originate from a public dataset [^5] Available at https://docs.prophesee.ai/stable/datasets.html. by Prophesee. For these experiments, we used the Event-to-Video method [^6] Available at https://bit.ly/Prophesee-E2VID. developed by Prophesee to generate a reference photo of the observed target (see our supplementary), and the EE3P method [10] with manual verification to estimate the GT frequency. All other sequences were captured as a part of this work and are publicly available. The final dataset contains sequences of period phenomena with frequencies ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM), see Fig. 3. Due to the non-constant frequencies of some periodic phenomena, we restricted our analysis to one-second segments of the event stream. This selection was based on the assumption that the change in frequency within such a time frame would be negligible and not significantly impact the accuracy of our measurements.
4.2 Measuring rotation speed
This section presents six experiments on measuring the rotation speed of objects.
4.2.1 Felt disc with a high-contrast line
For this experiment, we used a power drill to spin a white felt disc marked with a black line at $1200$ RPM (see our supplementary). Event data and tachometer readings were captured simultaneously.
All methods were expected to perform well because of the minimal noise and high contrast features present. Results confirmed this, with all methods achieving relative error lower than $0.001\%$ (see Tab. 4).
4.2.2 Fronto-parallel velcro disc
For this experiment, we used a more challenging target: a disc covered in uniform velcro material, where pattern recognition is difficult even for the human eye. A drill spun the disc at $1266$ RPM (see our supplementary).
Four out of six methods estimated the frequency despite the lack of clear features with relative error $<0.475\%$. The EB-ASM method achieved the worst accuracy (absolute error of $6228.9$ Hz, see Tab. 4). The Prophesee’s and our proposed method achieved the highest accuracy (relative error $<0.001$%), demonstrating robustness in this low-feature scenario.
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Target close-up (d) Aggregated events
Figure 4: Velcro disc with a non-frontal camera behind a glass sheet (see Sec. 4.2.3): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon. Positive events are represented by white colour, negative events are bright blue.
4.2.3 Velcro disc with a non-frontal camera behind a glass sheet
This experiment presented an even more challenging scenario. The velcro disc was spun at $1578$ RPM and captured at a $45\degree$ camera angle by the event camera through a glass sheet (see Fig. 4), simulating a possible industry application when the observed object is a rotating machine part and a transparent visor protects the camera and machine operators.
Interestingly, the Simple baseline method estimated the frequency very well ($0.38\%$ relative error). The output frequency of the FFT baseline method was roughly twice the actual frequency, likely due to the presence of centrosymmetric patterns on the velcro disc. The EB-ASM method failed to produce meaningful results. Even with the non-aligned camera axis resulting in elliptical object trajectories on the image plane, the method by Prophesee, Frequency-cam and our proposed methods produced accurate results ($<0.001$% relative error, see Tab. 4). The experiment confirms that capturing data through transparent materials does not compromise accuracy.
4.2.4 High-contrast dot
For this experiment, a sequence from Prophesee’s dataset of an orbiting dark dot was used, captured by an older event camera model with a lower resolution and an older sensor generation, resulting in the sequence being stored using an older file type (RAW EVT2777See https://docs.prophesee.ai/stable/data/encoding˙formats/index.html.). The Frequency-cam could not process this file encoding, hence the lack of measurements.
The Simple baseline method did not produce any meaningful results, and the FFT baseline method estimated a frequency approximately three times higher than the GT frequency (see Tab. 3). The Prophesee’s method achieved an absolute error of $0.1$ Hz, while EB-ASM and our proposed method achieved accurate frequency estimates ($<0.001$% relative error, see Tab. 4).
4.2.5 Hand fidget spinner
This experiment used event data from Prophesee’s dataset, capturing a spinning three-blade hand fidget spinner placed on a flat surface. Because of an older file encoding, the Frequency-cam method did not provide any measurements as its dependencies do not support reading the file.
Most of the methods were unable to provide any meaningful results. The FFT baseline method ($0.3$ Hz overestimation) and our proposed method were the only two that achieved good performance despite the object being centrosymmetric (see Tab. 4). Interestingly, the method by Prophesee produced a frequency approximately four times higher than the GT frequency despite the fidget spinner having only three blades. Our proposed method achieved the best overall performance ($<0.001$% relative error, see Tab. 4).
4.2.6 Whirling Pholcus phalangioides
The sequence used for this experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider. This unique defence mechanism involves the spider rapidly whirling its body while its legs remain at the silk [8], creating a challenging yet intriguing scenario for measuring the rotation speed of the body of the spider.
This experiment was characterised by significant noise, dim lighting conditions, and an unstable camera, which presented one of the most challenging sequences in the dataset. A second rotating pattern was visible, caused by the shadow of the spider (see our supplementary).
Most of the methods failed to provide accurate results. The FFT baseline method estimated a frequency four times higher than the GT frequency. The method by Prophesee and Frequency-cam methods produced estimates that deviated by $1$ Hz and $2.1$ Hz, respectively, from the GT frequency. Finally, our proposed method was the only one to achieve an accurate result with a relative error of $<0.001$% (see Tab. 4).
4.3 Measuring frequency of periodic light flashes
The following section presents experiments with frequency estimation methods for estimating high-frequency periodic flicker.
4.3.1 Flashing LED
A simple circuit with a light-emitting diode (see our supplementary) was used for this experiment. A software oscilloscope on the Raspberry Pi Pico controller [5] allowed precise control of the flashing frequency and duty cycle. The LED was set to flash exactly at $2000$ Hz with a $50\%$ duty cycle ($µ250\micro$s of turn-on time). Note that this frequency is substantially higher than those presented in other experiments in this work or related publications (see Sec. 2). The sequence used for this experiment was characterised by near-synchronous event generation across most pixels.
This experiment is the only other experiment (after the ’Felt disc with a high-contrast line’, Sec. 4.2.1) in which all methods achieved near-perfect results (relative error $<0.8\%$) with the Frequency-cam performing the worst (see Tab. 4) The EB-ASM and our proposed method achieved the lowest relative error ($<0.001$%).
4.3.2 Refreshing mobile phone screen
The sequence used for this experiment captured the flicker of a mobile phone screen that displayed a white colour at maximum brightness. A top-to-bottom pattern was visible when inspecting the event data, likely due to Pulse Width Modulation (PWM) for backlight control (see our supplementary). The GT frequency was estimated manually, using the EE3P method and confirmed by findings in a study on the display of this phone [^8] Available at https://bit.ly/notebookcheck-SamsungGalaxyS21UltraReview..
Although the Simple baseline method failed, all other methods achieved a relative error lower than $0.8\%$. The limitation of the proposed method is that it requires the aggregation of events to windows of length $t_{\text{quant}}$ (see Sec. 3.3). This restricts both the maximal measurable frequency and the theoretical limits of precision of the method. To illustrate, in this experiment, the ground truth period is $\frac{10^{6}}{240}=4166,\bar{6}$ microseconds. Therefore, the closest estimate of the period ($4200$ microseconds) results in an estimated frequency of $238.1$ Hz and a relative error of $0.792\%$. The Prophesee method and the FFT baseline method achieved the lowest relative errors ($<0.001$%, see Tab. 4).
4.4 Measuring vibration frequency
This section presents two experiments on measuring vibrations of a speaker diaphragm and a motor.
4.4.1 Speaker diaphragm
In this experiment, we captured and analysed a sequence capturing a vibrating diaphragm of a speaker with two large low-frequency drivers. An Android application [^9] Application available at https://bit.ly/Frequency-Sound-Generator-APK. precisely controlled the emitted sound (musical note $G_{2}$, $98$ Hz). The event camera captured the vibrating diaphragm of a speaker at approximately 30° angle. The most prominent features in the event stream originate from the vibrating edges of the diaphragm and the manufacturer’s logo located in its centre (see our supplementary).
The proposed method was the only method that achieved a relative error lower than $0.1\%$. The FFT baseline and Prophesee methods achieved relative error under $0.2\%$, followed by EB-ASM ($1.02\%$) and Frequency-cam ($2.8\%$). The Simple baseline method deviated the most from the reference frequency ($78.571\%$ relative error); see Tab. 4.
4.4.2 Vibrating motor
This experiment analysed a sequence from Prophesee’s dataset capturing a vibrating motor. The most reliable visual features were the horizontal cooling fins of the motor, which periodically moved in a vertical direction (see our supplementary).
The EB-ASM and the Simple baseline method deviated significantly from the GT frequency (relative error $>30\%$, see Tab. 4). Because of an older file encoding, the Frequency-cam method did not provide any measurements. The FFT baseline, Prophesee’s and our proposed method achieved a relative error lower than $0.001$%.
4.5 Measuring movement frequency
This section presents the application of frequency estimation methods for measuring the speed of moving objects. We present two experiments with a moving bicycle chain in which the frequency at which individual chain links pass is estimated.
4.5.1 Bike chain from side view
The sequence used for this experiment captured the bicycle chain from a close-up side view (see our supplementary).
The FFT baseline method correctly estimated the frequency in pixels that capture the upper and lower chain link sections, where events are periodically generated per chain link passage. However, the estimated frequency in most pixels was incorrect ($1$ Hz, see Tab. 3). The Prophesee’s method deviated by $\approx 18\%$ from the ground truth, being the second most accurate. Note that this is one of only four experiments in which the method by Prophesse achieved a relative error greater than $0.21\%$. Our proposed method achieved the best performance, being the only method to produce meaningful results (see Tab. 3).
4.5.2 Bike chain from top view
In the final experiment, we captured the same bike chain from a top view (see our supplementary).
The Frequency-cam method achieved the second-best result with a $1.3\%$ relative error, with the proposed method achieving the lowest absolute error of $0.1$ Hz. The Simple baseline, Prophesee’s, and EB-ASM methods showed significant deviations from the GT frequency (relative errors $34\%$$6111\%$).
4.6 Discussion
We analysed twelve sequences capturing various periodic phenomena using six different methods. The experiments where the target lacked high-contrast features were shown to be the most challenging. The worst performing methods were the Simple baseline method (see Sec. 3.1) and the EB-ASM method [2] achieving MRE of $4175.511\%$ and $6848.011\%$ respectively (see Tab. 3,4). The third lowest mean relative error across all experiments ($32.532\%$) was achieved by the Vibration Estimation method by Prophesee. The Frequency-cam method achieved the second lowest mean relative error (MRE) of $14.083\%$, which is significantly higher than the MRE achieved by the proposed method ($0.1\%$).
Note that the Frequency-cam method could not analyse three of the twelve sequences we used in this work, which originated from the public dataset by Prophesee due to an older file encoding type. All other methods were able to analyse all sequences.
Method GT Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam EE3P3D (ours)
1 Felt disc with $20.0$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$
high contrast line
2 Velcro disc (front) $21.1$ $24.2$ $21.0$ $\mathbf{21.1}$ $6250.0$ $21.2$ $\mathbf{21.1}$
3 Velcro disc (side) $26.3$ $26.4$ $53.0$ $\mathbf{26.3}$ $2439.0$ $\mathbf{26.3}$ $\mathbf{26.3}$
4 High contrast dot $19.6$ $7462.7$ $59.0$ $19.5$ $\mathbf{19.6}$ N/A $\mathbf{19.6}$
5 Fidget spinner $4.7$ $451.2$ $5.0$ $19.1$ $1$.11111.11111.11111.1 N/A $\mathbf{4.7}$
6 Whirling spider $3.2$ $35.1$ $13.0$ $4.2$ $298.6$ $5.3$ $\mathbf{3.2}$
7 Flashing LED $2000.0$ $1996.0$ $1999.4$ $2000.1$ $\mathbf{2000.0}$ $1984.1$ $\mathbf{2000.0}$
8 Refreshing mobile phone screen $240.0$ $2770.1$ $\mathbf{240.0}$ $\mathbf{240.0}$ $239.8$ $239.9$ $238.1$
9 Speaker diaphragm $98.0$ $21.0$ $98.1$ $97.8$ $99.0$ $95.3$ $\mathbf{98.0}$
10 Vibrating motor $40.0$ $52.2$ $\mathbf{40.0}$ $\mathbf{40.0}$ $370.4$ N/A $\mathbf{40.0}$
11 Bike chain (side) $28.7$ $123.8$ $1.0$ $33.9$ $1$.11111.11111.11111.1 $44.7$ $\mathbf{28.7}$
12 Bike chain (top) $23.0$ $51.6$ $22.0$ $30.8$ $1428.6$ $22.7$ $\mathbf{23.1}$
Table 3: Frequency estimates (Hz) on the 12 sequence dataset. Methods: Simple baseline (Sec. 3.1), FFT baseline (Sec. 3.2), Prophesse Vibration Estimation [14], EB-ASM [2], Frequency-cam [13] and EE3P3D (Sec. 3.3).
Method Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam EE3P3D (ours)
Felt disc with high contrast line $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$
Velcro disc (front) $14.6$ $0.5$ $<\epsilon$ $29520.9$ $0.5$ $<\epsilon$
Velcro disc (side) $0.4$ $101.5$ $<\epsilon$ $9173.8$ $<\epsilon$ $<\epsilon$
High contrast dot $37975.0$ $201.0$ $0.5$ $<\epsilon$ N/A $<\epsilon$
Fidget spinner $9500.0$ $6.3$ $306.3$ $23540.4$ N/A $<\epsilon$
Whirling spider $996.9$ $306.3$ $31.3$ $9231.3$ $65.6$ $<\epsilon$
Flashing LED $0.2$ $0.03$ $0.005$ $<\epsilon$ $0.7$ $<\epsilon$
Refreshing mobile phone screen $1054.2$ $<\epsilon$ $<\epsilon$ $0.08$ $0.04$ $0.8$
Speaker diaphragm $78.6$ $0.1$ $0.2$ $1.0$ $2.8$ $<\epsilon$
Vibrating motor $30.5$ $<\epsilon$ $<\epsilon$ $826.0$ N/A $<\epsilon$
Bike chain (side) $331.4$ $96.5$ $18.1$ $3771.4$ $55.8$ $<\epsilon$
Bike chain (top) $124.3$ $4.3$ $33.9$ $6111.3$ $1.3$ $0.4$
Mean relative error $4175.5$ $59.7$ $32.5$ $6848.0$ $14.0$ $0.1$
Max relative error $37975.0$ $306.2$ $306.3$ $29520.9$ $65.6$ $0.8$
Table 4: Relative errors (%) on the 12 sequence dataset. Methods: Simple baseline (Sec. 3.1), FFT baseline (Sec. 3.2), Prophesse Vibration Estimation [14], EB-ASM [2], Frequency-cam [13] and EE3P3D (Sec. 3.3), $\epsilon=10^{-3}$.
5 Conclusion
In this work, a novel contactless method for measuring the frequency of periodic phenomena is presented. The method makes only a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream.
The proposed method, EE3P3D, is evaluated on twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), with targets lacking prominent features, being behind transparent materials, and having centrosymmetric shapes, achieving the mean relative error of $0.1\%$. The method works in non-frontal camera placement, measuring frequencies as high as $2000$ Hz (equivalent to 120 000120000120\,000120 000 RPM) even when a significant amount of noise is present.
The dataset and codes are publicly available on GitHub.
Appendix 0.A Experimental Setup
For the ’Felt disc with a high-contrast line’ experiment (see Sec. 4.2.1), we captured a power drill spinning a white felt disc marked with a black line, see Fig. 0.A.5.
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.5: Felt disc with a high-contrast line (see Sec. 4.2.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($655\times 655$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
For the ’Fronto-parallel velcro disc’ experiment (see Sec. 4.2.2), a drill spun the disc at $1266$ RPM with the event camera facing it directly, while a laser tachometer measured speed from the side of the disc where a reflective sticker was attached (Fig. 0.A.6a).
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) RoI
Figure 0.A.6: Fronto-parallel velcro disc (see Sec. 4.2.2): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($713\times 713$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
In the next experiment (’Velcro disc (side)’, see Sec. 4.2.3), the velcro disc was captured by the event camera through a glass sheet at $45\degree$ camera angle (see Fig. 4a).
For the following experiment (’High contrast dot’), a sequence from Prophesee’s dataset was used, capturing an orbiting dark dot (see Sec. 4.2.4). The reference photo was generated using Prophesee’s Event to Video method (see Fig. 0.A.7a).
[Uncaptioned image] [Uncaptioned image]
(a) Target visualisation (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.7: High-contrast dot (see Sec. 4.2.4): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($263\times 263$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
This experiment (see Sec. 4.2.5) used event data from Prophesee’s dataset, capturing a rotating three-blade hand fidget spinner on a flat surface, see Fig. 0.A.8.
[Uncaptioned image] [Uncaptioned image]
(a) Target visualisation (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.8: Hand fidget spinner (see Sec. 4.2.5): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($239\times 239$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for the next experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider (see Sec. 4.2.6). A second rotating pattern is visible, caused by the shadow of the spider (see Fig. 0.A.9d,e).
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(a) Physical setup                 (b) Target close-up (c) Region of Interest
[Uncaptioned image] [Uncaptioned image]
(d) Aggregated events (e) XYTime visualisation
Figure 0.A.9: Whirling Pholcus phalangioides (see Sec. 4.2.6): (a) physical setup; (b) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($221\times 221$px) shown in (c); (e) events from a 250-millisecond window visualised in spatio-temporal space. Positive events are represented by white colour, and negative events are bright blue.
For the LED experiment (see Sec. 4.3.1), a circuit with a light-emitting diode (see Fig. 0.A.10a) was used.
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.10: Flashing LED (see Sec. 4.3.1): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($126\times 126$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for the ’Refreshing mobile phone screen’ experiment (see Sec. 4.3.2) captured the flicker of a mobile phone screen displaying a white colour at maximum brightness. A moving top-to-bottom pattern was visible when inspecting the event data (see Fig. 0.A.11b).
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.11: Refreshing mobile phone screen (see Sec. 4.3.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within hardware-level region of interest and a time window equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($54\times 54$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
In the following experiment (see Sec. 4.4.1), we captured a vibrating diaphragm of a speaker with two large low-frequency drivers, see Fig. 0.A.12a.
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.12: Speaker diaphragm (see Sec. 4.4.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($631\times 631$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
For the ’Vibrating motor’ experiment (see Sec. 4.4.2) we used a sequence from Prophesee’s dataset capturing a motor vibrating vertically at $40$ Hz; see Fig. 0.A.13.
[Uncaptioned image] [Uncaptioned image]
(a) Target visualisation (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.13: Vibrating motor (see Sec. 4.4.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($148\times 148$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for this experiment (see Sec. 4.5.1) captured a bicycle chain from a close-up side view (see Fig. 0.A.14a).
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.14: Bike chain from side view (see Sec. 4.5.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within a hardware-level region of interest and a time window equal to the period of the observed phenomenon, highlighted region of interest ($204\times 204$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence for the final experiment (see Sec. 4.5.2) captured the same bike chain from a top view perspective (see Fig. 0.A.15a).
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.15: Bike chain from a top view (see Sec. 4.5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of lenght equal to the period of the observed phenomenon, highlighted region of interest ($180\times 180$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.

### v2

@@
-EE3P3D: Event-based Estimation of Periodic Phenomena Frequency using 3D Correlation
+EEPPR: Event-based Estimation of Periodic Phenomena Rate using Correlation in 3D
 Abstract
-We present a novel method for measuring the frequency of periodic phenomena, e.g., rotation, flicker and vibration, by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution.
-The approach assumes that for a periodic phenomenon, a highly similar set of events is generated within a specific spatio-temporal window at a time difference corresponding to the phenomenon’s period.
-The sets of similar events are detected by 3D spatio-temporal correlation in the event stream space.
-The proposed method, EE3P3D, is evaluated on a dataset of $12$ sequences of periodic phenomena, i.e.
+We present a novel method for measuring the period of phenomena like rotation, flicker and vibration, by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution.
+The approach assumes that for a periodic phenomenon, a highly similar set of events is generated within a spatio-temporal window at a time difference corresponding to its period.
+The sets of similar events are detected by a correlation in the spatio-temporal event stream space.
+The proposed method, EEPPR, is evaluated on a dataset of $12$ sequences of periodic phenomena, i.e.
 flashing light and vibration, and periodic motion, e.g., rotation, ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM).
-EE3P3D significantly outperforms published methods on this dataset, achieving the mean relative error of $0.1\%$.
+EEPPR significantly outperforms published methods on this dataset, achieving the mean relative error of $0.1\%$.
+The dataset and codes are available on GitHub.
 Keywords:
-Event camera Frequency estimation Periodic phenomena Frequency of periodic phenomena.
+Event camera, Frequency estimation, Periodic phenomena, Frequency of periodic phenomena.
 1 Introduction
@@
 Precise quantification of rotational speed, for instance, finds applications in (i) monitoring rotating machinery components for performance evaluation and quality control [4], (ii) ensuring flight stability and manoeuvrability in drones [15], (iii) analysing sports equipment like ball tracking [9, 7], and (iv) optimising energy production in wind turbines [16].
-Traditional methods for measuring periodic phenomena often involve contact-based devices such as tachometers and rotary encoders.
+Traditional methods for measuring periodic phenomena often involve contact-based devices such as mechanical tachometers and rotary encoders.
 These approaches, while established, have inherent limitations: (i) physical contact with the target object can influence its movement and introduce measurement inaccuracies, (ii) contact methods may not be feasible in scenarios with delicate objects, confined spaces, or situations where interfering with the target significantly affects its movement.
@@
 Although contactless event-based methods exist, they are either handcrafted for a specific type of periodic phenomena [1] or require user supervision [13, 14].
-Figure 1: EE3P3D: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a 3D correlation of the template with the event stream is computed, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows.
+Figure 1: EEPPR method: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a 3D correlation of the template with the event stream is computed, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows.
 In this paper, we propose a non-contact method for measuring the properties of any periodic phenomena using an event camera.
-The input of our method is the event stream, and the output is a single scalar – the frequency of the periodic phenomena.
-Our approach assumes that similar sets of pixel activations of an event camera (events) will occur at periodic time intervals in any sequence capturing a periodic phenomenon.
-The periodic intervals correspond to the target frequency.
-We evaluated our method on a dataset of twelve sequences of four types of periodic phenomena: (i) light flickering, (ii) object vibration, (iii) rotational speed, and (iv) periodic movement and achieved a mean relative error of $0.1\%$, which is 140$\approx$ 140$\times$ 140 × lower than the mean relative error achieved by the best available published method (see Tab.
-4).
+The input of our method is the event stream, and the output is a single scalar – the rate of the periodic phenomena.
+The method operates remotely, without requiring any modification of the observed object (e.g.
+a marker).
+The approach assumes that similar sets of pixel activations of an event camera (events) will occur at regular time intervals in any sequence capturing a periodic phenomenon.
+These intervals correspond to the period.
+The method was evaluated on a new dataset of twelve sequences of four types of periodic phenomena: (i) flickering light, (ii) object vibration, (iii) rotation, and (iv) repetitive translational movement.
+A mean relative error of $0.1\%$ was achieved, which is 140$\approx$ 140$\times$ 140 × lower than the mean relative error of the best available published method (see Tab.
+3).
+The point estimate for the period lengths is obtained by calculating a median in a spatial window for robustness.
+For phenomena where the time of repetition varies, the method provides the time of each period which can be further analysed.
 2 Related work
@@
 Electrostatic sensors offer a contactless approach, detecting changes in the electromagnetic field around a rotating object (equipped with a bearing).
-The rotational speed is the frequency of these changes [11].
+The rotational speed is the rate of these changes [11].
 Optical encoder tachometers use a photoelectric sensor to detect light passing through a rotating disc with patterned segments.
@@
 Although their approach is unknown, we present the results of their method in our experiments.
-3 Methods
-In this section, we describe three proposed methods. The Simple (Sec. 3.1) and FFT (Sec. 3.2) methods serve as baselines, and the proposed method is presented in Sec. 3.3.
-3.1 A Simple baseline
-A baseline method was implemented to test the hypothesis that separate per-pixel events analysis is insufficient for accurate frequency estimation of complex periodic phenomena.
+Window size (px) $30\times 30$ $45\times 45$ $60\times 60$ $75\times 75$
+Avg relative error $37.78$ $\mathbf{19.34}$ $42.96$ $487.98$
+Max relative error $\mathbf{3225.00}$ $3362.50$ $7712.50$ $22737.50$
+Table 1: EEPPR, average and maximum relative errors (%) as functions of window size on the dataset 5.1.
+Template ev count $100-1400$ $1500-3000$
+Avg relative error $308.46$ $\mathbf{5.75}$
+Max relative error $22737.50$ $\mathbf{269.36}$
+Table 2: EEPPR, dependence on template event count (see Sec. 4) on the dataset 5.1.
+3 Baselines
+A Simple baseline
+was implemented to test the hypothesis that separate per-pixel events analysis is insufficient for accurate estimation of rate of complex periodic phenomena.
 The event camera produces a sequence of tuples $(x,y,p,t)$, where $x,y$ are spatial coordinates, $p$ is event polarity (positive – $1$ or negative – 00), and $t$ is a timestamp in microseconds of the event creation.
@@
   3. 3.
-    Frequency prediction – the final predicted frequency for each window is derived from the estimated period $T$ in microseconds using the following equation $\nu(T)={10^{6}}/{T}$, where $\nu(T)$ returns frequency in Hertz (Hz).
-3.2 A FFT baseline
-The FFT baseline method estimates per-pixel frequency in four steps:
+    Rate prediction – the final predicted rate for each window is derived from the estimated period $T$ in microseconds using the following equation $\nu(T)={10^{6}}/{T}$, where $\nu(T)$ returns rate in Hertz (Hz).
+An FFT baseline
+estimates per-pixel rate in four steps:
   1. 1.
-    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform with the efficient Fast Fourier Transform (FFT) algorithm [3] with the timestamp of event creation represented as a Dirac pulse.
+    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform with the Fast Fourier Transform (FFT)  [3] with the event creation time represented by a Dirac pulse.
   2. 2.
@@
   3. 3.
-    Median filtering – a 2D median filter using a local window size of $3$ is applied to the 2D array of estimated per-pixel frequencies to mitigate outliers.
+    Median filtering – in a local 3x3 window is applied to the 2D array of estimated per-pixel rates to mitigate outliers.
   4. 4.
-    Results aggregation – the mode of detected frequencies is returned.
-3.3 EE3P3D - The proposed method
-Broadly speaking, the event stream is quantized and 3D correlation responses analysed to detect periodic similarities along the time axis of the event data.
-An overview of the method is shown in Fig.
-1.
-We identify the period of a spatio-temporal pattern in the event stream in four steps:
+    Results aggregation – the mode of detected rate is returned.
+4 EEPPR – The proposed method
+Our method (see Fig. 1) detects periodic similarities along the time axis of event data in four steps:
   1. 1.
@@
   3. 3.
-    3D correlation – a 3D correlation of the template $\mathbf{T}_{i}$ is computed with the area $\mathbf{A}_{i}$ (see Fig.
+    Correlation in 3D – a correlation in 3D of the template $\mathbf{T}_{i}$ is computed with the area $\mathbf{A}_{i}$ (see Fig.
 2).
@@
 5 Experiments
-In this section, we discuss the event camera used and the selection of parameters for methods presented in the previous section, followed by experiments measuring rotation speed, frequency of periodic light flashes, vibrations, and moving objects.
-In our experiments, we used the Prophesee EVK4 HD event camera with $1280\times 720$ resolution, capturing up to $1066$ million events per second [6] under light conditions ranging from $0.08$ up to 100 000100000100\,000100 000 lux [^2] Spec sheet available at https://bit.ly/Sony-Prophesee-IMX636-IMX637-specsheet.
+In the experiments, we used the Prophesee EVK4 HD event camera with $1280\times 720$ resolution, capturing up to $1066$ million events per second [6] under light conditions ranging from $0.08$ up to 100 000100000100\,000100 000 lux [^2] Spec sheet available at https://bit.ly/Sony-Prophesee-IMX636-IMX637-specsheet.
 with microsecond time resolution (equivalent to 10 0001000010\,00010 000 frames per second).
@@
 We left the remaining biases at their default settings.
+Since the periodic phenomena in our dataset are stationary, and there is no ego motion of the camera, we perform 3D correlation only along the time axis – “correlation in 3D”.
+However, in cases where the actual 3D correlation would be beneficial, it can be applied without any issues.
 We evaluated the accuracy of the Simple baseline method (see Sec.
-3.1) when using window sizes $W$ of $30\times 30$ px, $45\times 45$ px, $60\times 60$ px and $75\times 75$ px and achieved the lowest mean relative error (MRE) by setting $W$ to $45\times 45$ px.
+3) when using window sizes $W$ of $30\times 30$ px, $45\times 45$ px, $60\times 60$ px and $75\times 75$ px and achieved the lowest mean relative error (MRE) by setting $W$ to $45\times 45$ px.
 Further increasing or decreasing window size or using mode instead of the median in the temporal or spatial analysis steps of this method negatively impacted the accuracy.
-The FFT baseline method (see Sec. 3.2) achieved lower mean relative error when using timestamps of events with negative polarity (MRE of $60\%$ as opposed to $67.467\%$).
+The FFT baseline method (see Sec.
+3) achieved lower mean relative error when using timestamps of events with negative polarity (MRE of $60\%$ as opposed to $67.467\%$).
 We evaluated all different configurations of parameters (window sizes $W$ of $30$, $45$, $60$ and $75$ and counts of template events $N$ of $100$, $200$, …, $3000$) for the proposed method (see Sec.
-3.3).
+4).
 In all our experiments, the length of the quantisation window $t_{\text{quant}}$ is set to $100$ microseconds.
 When using window size $W$ set to $45\times 45$ px the method achieves the lowest MRE by a substantial margin and a low maximal relative error compared to results when using other window sizes (see Tab.
-1).
+2).
 We noticed a significant improvement in accuracy when using counts of template events $N$ higher than $1500$ (see Tab.
 2).
-The best accuracy was achieved using the window size $W$ set to $45\times 45$ px and the template event count $N$ set to $1800$ events.
+Using 3-fold cross-validation on the introduced dataset, the best accuracy is achieved using the window size $W$ set to $45\times 45$ px and the template event count $N$ set to $1800$ events.
 These parameters are proposed as default parameters of the method.
-To objectively compare the results of the proposed method, we use the mentioned baseline methods (see Sec.
-3.1, 3.2), a vibration estimation method [14] developed by Prophesee, manufacturer of the used event camera, the open-source Frequency-cam method [13] and re-implemented EB-ASM method [2].
-As the output of Frequency-cam is in the form of colourmaps, we modified it to instead output a reading for each pixel that would be used for colourmap creation and computed their median.
+We compare the results of the proposed method both with the baseline methods (see Sec.
+3), and published methods – a vibration estimation method [14] developed by Prophesee, manufacturer of the used event camera, the open-source Frequency-cam method [13] and EB-ASM method [1] which we re-implemented.
+As the output of Frequency-cam is in the form of colourmaps, we modified it to output a reading for each pixel instead and computed their median.
 Since every method produces a different amount of measurements per one second of input data, we calculate the median of measurements to obtain a single value for a comparison with other methods.
-Window size (px) $30\times 30$ $45\times 45$ $60\times 60$ $75\times 75$
-Avg rel error $37.78$ $\mathbf{19.34}$ $42.96$ $487.98$
-Max rel error $\mathbf{3225.00}$ $3362.50$ $7712.50$ $22737.50$
-Table 1: EE3P3D - dependence on window size. Average and maximum relative errors (%) on the 12 sequence dataset.
-Template event count $100-1400$ $1500-3000$
-Average relative error $308.46$ $\mathbf{5.75}$
-Maximum relative error $22737.50$ $\mathbf{269.36}$
-Table 2: EE3P3D - dependence on template event count (see Sec. 3.3). Average and maximum relative errors (%) on the $12$ sequence dataset.
+We followed parameter selection guidelines provided by the authors of all examined methods [14, 13, 1].
 5.1 Dataset
@@
 We averaged all tachometer readings captured in one second ($3$ to $5$) to obtain a single reference value for comparison with the results of all methods.
-Figure 3: Ground-truth frequencies of experiments in order of appearance in this work.
+Figure 3: Ground-truth rates of experiments in order of appearance in this work.
 However, acquiring GT data this way was not feasible for all experiments.
-In such cases, the EE3P method [10] was used to analyse the data multiple times with varying parameter configurations to filter potential measurement outliers and obtain a GT frequency estimate as close as possible to the actual value.
-The frequency was then verified by manually examining the event data stream.
+In such cases, the EE3P method [10] was used to analyse the data multiple times with varying parameter configurations to filter potential measurement outliers and obtain a GT rate estimate as close as possible to the actual value.
+The rate was then verified by manually examining the event data stream.
 Event data for three experiments in this work originate from a public dataset [^5] Available at https://docs.prophesee.ai/stable/datasets.html.
 by Prophesee.
-For these experiments, we used the Event-to-Video method [^6] Available at https://bit.ly/Prophesee-E2VID.
-developed by Prophesee to generate a reference photo of the observed target (see our supplementary), and the EE3P method [10] with manual verification to estimate the GT frequency.
+For these experiments, we used the EE3P method [10] with manual verification to estimate the GT rate.
 All other sequences were captured as a part of this work and are publicly available.
-The final dataset contains sequences of period phenomena with frequencies ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM), see Fig.
+The final dataset contains sequences of period phenomena with rates ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM), see Fig.
 3.
-Due to the non-constant frequencies of some periodic phenomena, we restricted our analysis to one-second segments of the event stream.
-This selection was based on the assumption that the change in frequency within such a time frame would be negligible and not significantly impact the accuracy of our measurements.
+Due to the non-constant rates of some periodic phenomena, we restricted our analysis to one-second segments of the event stream.
+This selection was based on the assumption that the change in rate within such a time frame would be negligible and not significantly impact the accuracy of our measurements.
 5.2 Measuring rotation speed
-This section presents six experiments on measuring the rotation speed of objects.
-4.2.1 Felt disc with a high-contrast line
-For this experiment, we used a power drill to spin a white felt disc marked with a black line at $1200$ RPM (see our supplementary). Event data and tachometer readings were captured simultaneously.
+Felt disc with a high-contrast line
+For this experiment, we used a power drill to spin a white felt disc marked with a black line at $1200$ RPM.
+Event data and tachometer readings were captured simultaneously.
 All methods were expected to perform well because of the minimal noise and high contrast features present.
 Results confirmed this, with all methods achieving relative error lower than $0.001\%$ (see Tab.
-4).
-4.2.2 Fronto-parallel velcro disc
+3).
+Fronto-parallel velcro disc
 For this experiment, we used a more challenging target: a disc covered in uniform velcro material, where pattern recognition is difficult even for the human eye.
-A drill spun the disc at $1266$ RPM (see our supplementary).
-Four out of six methods estimated the frequency despite the lack of clear features with relative error $<0.475\%$.
+A drill spun the disc at $1266$ RPM.
+Four out of six methods estimated the rate despite the lack of clear features with relative error $<0.475\%$.
 The EB-ASM method achieved the worst accuracy (absolute error of $6228.9$ Hz, see Tab.
-4).
+3).
 The Prophesee’s and our proposed method achieved the highest accuracy (relative error $<0.001$%), demonstrating robustness in this low-feature scenario.
@@
 Figure 4: Velcro disc with a non-frontal camera behind a glass sheet (see Sec.
-4.2.3): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon.
+5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon.
 Positive events are represented by white colour, negative events are bright blue.
-4.2.3 Velcro disc with a non-frontal camera behind a glass sheet
+Velcro disc with a non-frontal camera behind a glass sheet
 This experiment presented an even more challenging scenario.
@@
 4), simulating a possible industry application when the observed object is a rotating machine part and a transparent visor protects the camera and machine operators.
-Interestingly, the Simple baseline method estimated the frequency very well ($0.38\%$ relative error).
-The output frequency of the FFT baseline method was roughly twice the actual frequency, likely due to the presence of centrosymmetric patterns on the velcro disc.
+Interestingly, the Simple baseline method estimated the rate very well ($0.38\%$ relative error).
+The output frequency of the FFT baseline method was roughly twice the actual rate, likely due to the presence of centrosymmetric patterns on the velcro disc.
 The EB-ASM method failed to produce meaningful results.
 Even with the non-aligned camera axis resulting in elliptical object trajectories on the image plane, the method by Prophesee, Frequency-cam and our proposed methods produced accurate results ($<0.001$% relative error, see Tab.
-4).
+3).
 The experiment confirms that capturing data through transparent materials does not compromise accuracy.
-4.2.4 High-contrast dot
-For this experiment, a sequence from Prophesee’s dataset of an orbiting dark dot was used, captured by an older event camera model with a lower resolution and an older sensor generation, resulting in the sequence being stored using an older file type (RAW EVT2777See https://docs.prophesee.ai/stable/data/encoding˙formats/index.html.).
+High-contrast dot
+A sequence from Prophesee’s dataset of an orbiting dark dot was used, captured by an older event camera model with a lower resolution and an older sensor generation, resulting in the sequence being stored using an older file type (RAW EVT2666See https://docs.prophesee.ai/stable/data/encoding˙formats/index.html.).
 The Frequency-cam could not process this file encoding, hence the lack of measurements.
-The Simple baseline method did not produce any meaningful results, and the FFT baseline method estimated a frequency approximately three times higher than the GT frequency (see Tab.
+The Simple baseline method did not produce any meaningful results, and the FFT baseline method estimated a frequency approximately three times higher than the GT rate.
+The Prophesee’s method achieved an absolute error of $0.1$ Hz, while EB-ASM and our proposed method achieved accurate rate estimates ($<0.001$% relative error, see Tab.
 3).
-The Prophesee’s method achieved an absolute error of $0.1$ Hz, while EB-ASM and our proposed method achieved accurate frequency estimates ($<0.001$% relative error, see Tab.
-4).
-4.2.5 Hand fidget spinner
+Hand fidget spinner
 This experiment used event data from Prophesee’s dataset, capturing a spinning three-blade hand fidget spinner placed on a flat surface.
@@
 Most of the methods were unable to provide any meaningful results.
-The FFT baseline method ($0.3$ Hz overestimation) and our proposed method were the only two that achieved good performance despite the object being centrosymmetric (see Tab.
-4).
-Interestingly, the method by Prophesee produced a frequency approximately four times higher than the GT frequency despite the fidget spinner having only three blades.
+The FFT baseline method ($0.3$ Hz overestimation) and our proposed method were the only two that achieved good performance despite the object being centrosymmetric.
+Interestingly, the method by Prophesee produced a rate approximately four times higher than the GT rate despite the fidget spinner having only three blades.
 Our proposed method achieved the best overall performance ($<0.001$% relative error, see Tab.
-4).
-4.2.6 Whirling Pholcus phalangioides
+3).
+Whirling Pholcus phalangioides
 The sequence used for this experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider.
@@
 This experiment was characterised by significant noise, dim lighting conditions, and an unstable camera, which presented one of the most challenging sequences in the dataset.
-A second rotating pattern was visible, caused by the shadow of the spider (see our supplementary).
+A second rotating pattern was visible, caused by the shadow of the spider.
 Most of the methods failed to provide accurate results.
-The FFT baseline method estimated a frequency four times higher than the GT frequency.
-The method by Prophesee and Frequency-cam methods produced estimates that deviated by $1$ Hz and $2.1$ Hz, respectively, from the GT frequency.
+The FFT baseline method estimated a frequency four times higher than the GT rate.
+The method by Prophesee and Frequency-cam methods produced estimates that deviated by $1$ Hz and $2.1$ Hz, respectively, from the GT rate.
 Finally, our proposed method was the only one to achieve an accurate result with a relative error of $<0.001$% (see Tab.
-4).
+3).
 5.3 Measuring frequency of periodic light flashes
-The following section presents experiments with frequency estimation methods for estimating high-frequency periodic flicker.
-4.3.1 Flashing LED
-A simple circuit with a light-emitting diode (see our supplementary) was used for this experiment.
+Flashing LED
+A simple circuit with a light-emitting diode was used for this experiment.
 A software oscilloscope on the Raspberry Pi Pico controller [5] allowed precise control of the flashing frequency and duty cycle.
@@
 This experiment is the only other experiment (after the ’Felt disc with a high-contrast line’, Sec.
-4.2.1) in which all methods achieved near-perfect results (relative error $<0.8\%$) with the Frequency-cam performing the worst (see Tab.
-4) The EB-ASM and our proposed method achieved the lowest relative error ($<0.001$%).
-4.3.2 Refreshing mobile phone screen
+5.2) in which all methods achieved near-perfect results (relative error $<0.8\%$) with the Frequency-cam performing the worst (see Tab.
+3) The EB-ASM and our proposed method achieved the lowest relative error ($<0.001$%).
+Refreshing screen
 The sequence used for this experiment captured the flicker of a mobile phone screen that displayed a white colour at maximum brightness.
-A top-to-bottom pattern was visible when inspecting the event data, likely due to Pulse Width Modulation (PWM) for backlight control (see our supplementary).
-The GT frequency was estimated manually, using the EE3P method and confirmed by findings in a study on the display of this phone [^8] Available at https://bit.ly/notebookcheck-SamsungGalaxyS21UltraReview..
+A top-to-bottom pattern was visible when inspecting the event data, likely due to Pulse Width Modulation (PWM) for backlight control.
+The GT rate was estimated manually, using the EE3P method and confirmed by findings in a study on the display of this phone [^7] Available at https://bit.ly/notebookcheck-SamsungGalaxyS21UltraReview..
 Although the Simple baseline method failed, all other methods achieved a relative error lower than $0.8\%$.
 The limitation of the proposed method is that it requires the aggregation of events to windows of length $t_{\text{quant}}$ (see Sec.
-3.3).
+4).
 This restricts both the maximal measurable frequency and the theoretical limits of precision of the method.
 To illustrate, in this experiment, the ground truth period is $\frac{10^{6}}{240}=4166,\bar{6}$ microseconds.
-Therefore, the closest estimate of the period ($4200$ microseconds) results in an estimated frequency of $238.1$ Hz and a relative error of $0.792\%$.
+Therefore, the closest estimate of the period ($4200$ microseconds) results in an estimated rate of $238.1$ Hz and a relative error of $0.792\%$.
 The Prophesee method and the FFT baseline method achieved the lowest relative errors ($<0.001$%, see Tab.
+3).
+Frequency Estimates (Hz) Relative Errors (%)
+Method GT Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam Ours Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam Ours
+1 Felt disc with high contrast line 20.0 20.0 20.0 20.0 20.0 20.0 20.0 $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$
+2 Velcro disc (front) 21.1 24.2 21.0 21.1 6250.0 21.2 21.1 14.6 0.5 $<\epsilon$ 29520.9 0.5 $<\epsilon$
+3 Velcro disc (side) 26.3 26.4 53.0 26.3 2439.0 26.3 26.3 0.4 101.5 $<\epsilon$ 9173.8 $<\epsilon$ $<\epsilon$
+4 High contrast dot 19.6 7462.7 59.0 19.5 19.6 N/A 19.6 37975.0 201.0 0.5 $<\epsilon$ N/A $<\epsilon$
+5 Fidget spinner 4.7 451.2 5.0 19.1 $1$.1 N/A 4.7 9500.0 6.3 306.3 23540.4 N/A $<\epsilon$
+6 Whirling spider 3.2 35.1 13.0 4.2 298.6 5.3 3.2 996.9 306.3 31.3 9231.3 65.6 $<\epsilon$
+7 Flashing LED 2000.0 1996.0 1999.4 2000.1 2000.0 1984.1 2000.0 0.2 0.03 0.005 $<\epsilon$ 0.7 $<\epsilon$
+8 Refreshing screen 240.0 2770.1 240.0 240.0 239.8 239.9 238.1 1054.2 $<\epsilon$ $<\epsilon$ 0.08 0.04 0.8
+9 Speaker diaphragm 98.0 21.0 98.1 97.8 99.0 95.3 98.0 78.6 0.1 0.2 1.0 2.8 $<\epsilon$
+10 Vibrating motor 40.0 52.2 40.0 40.0 370.4 N/A 40.0 30.5 $<\epsilon$ $<\epsilon$ 826.0 N/A $<\epsilon$
+11 Bike chain (side) 28.7 123.8 1.0 33.9 $1$.1 44.7 28.7 331.4 96.5 18.1 3771.4 55.8 $<\epsilon$
+12 Bike chain (top) 23.0 51.6 22.0 30.8 1428.6 22.7 23.1 124.3 4.3 33.9 6111.3 1.3 0.4
+Table 3: Frequency estimates (Hz) and relative errors (%) on the 12 sequence dataset.
+Methods: Simple baseline (Sec.
+3), FFT baseline (Sec.
+3), Prophesee Vibration Estimation [14], EB-ASM [1], Frequency-cam [13] and EEPPR (Sec.
 4).
-4.4 Measuring vibration frequency
-This section presents two experiments on measuring vibrations of a speaker diaphragm and a motor.
-4.4.1 Speaker diaphragm
+5.4 Measuring vibration frequency of a speaker and a motor
+Speaker diaphragm
 In this experiment, we captured and analysed a sequence capturing a vibrating diaphragm of a speaker with two large low-frequency drivers.
@@
 The event camera captured the vibrating diaphragm of a speaker at approximately 30° angle.
-The most prominent features in the event stream originate from the vibrating edges of the diaphragm and the manufacturer’s logo located in its centre (see our supplementary).
+The most prominent features in the event stream originate from the vibrating edges of the diaphragm and the manufacturer’s logo located in its centre.
 The proposed method was the only method that achieved a relative error lower than $0.1\%$.
@@
 The Simple baseline method deviated the most from the reference frequency ($78.571\%$ relative error); see Tab.
-4.
-4.4.2 Vibrating motor
+3.
+Vibrating motor
 This experiment analysed a sequence from Prophesee’s dataset capturing a vibrating motor.
-The most reliable visual features were the horizontal cooling fins of the motor, which periodically moved in a vertical direction (see our supplementary).
+The most reliable visual features were the horizontal cooling fins of the motor, which periodically moved in a vertical direction.
 The EB-ASM and the Simple baseline method deviated significantly from the GT frequency (relative error $>30\%$, see Tab.
-4).
+3).
 Because of an older file encoding, the Frequency-cam method did not provide any measurements.
@@
 5.5 Measuring movement frequency
-This section presents the application of frequency estimation methods for measuring the speed of moving objects.
-We present two experiments with a moving bicycle chain in which the frequency at which individual chain links pass is estimated.
-4.5.1 Bike chain from side view
-The sequence used for this experiment captured the bicycle chain from a close-up side view (see our supplementary).
+Bike chain from side view
+The sequence used for this experiment captured the bicycle chain from a close-up side view.
 The FFT baseline method correctly estimated the frequency in pixels that capture the upper and lower chain link sections, where events are periodically generated per chain link passage.
@@
 3).
-The Prophesee’s method deviated by $\approx 18\%$ from the ground truth, being the second most accurate.
-Note that this is one of only four experiments in which the method by Prophesse achieved a relative error greater than $0.21\%$.
-Our proposed method achieved the best performance, being the only method to produce meaningful results (see Tab.
-3).
-4.5.2 Bike chain from top view
-In the final experiment, we captured the same bike chain from a top view (see our supplementary).
+The Prophesee method deviated by $\approx 18\%$ from the ground truth, being the second most accurate; this is one of only four experiments where the method by Prophesse had a relative error greater than $0.21\%$.
+EEPPR achieved the best performance, being the only method to produce meaningful results.
+Bike chain from top view
+In the final experiment, we captured the same bike chain from a top view.
 The Frequency-cam method achieved the second-best result with a $1.3\%$ relative error, with the proposed method achieving the lowest absolute error of $0.1$ Hz.
@@
 The experiments where the target lacked high-contrast features were shown to be the most challenging.
-The worst performing methods were the Simple baseline method (see Sec.
-3.1) and the EB-ASM method [2] achieving MRE of $4175.511\%$ and $6848.011\%$ respectively (see Tab.
-3,4).
+The worst performing methods were the Simple baseline method and the EB-ASM method [1] achieving MRE of $4175.511\%$ and $6848.011\%$ respectively (see Tab.
+3).
 The third lowest mean relative error across all experiments ($32.532\%$) was achieved by the Vibration Estimation method by Prophesee.
@@
 All other methods were able to analyse all sequences.
-Method GT Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam EE3P3D (ours)
-1 Felt disc with $20.0$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$ $\mathbf{20.0}$
-high contrast line
-2 Velcro disc (front) $21.1$ $24.2$ $21.0$ $\mathbf{21.1}$ $6250.0$ $21.2$ $\mathbf{21.1}$
-3 Velcro disc (side) $26.3$ $26.4$ $53.0$ $\mathbf{26.3}$ $2439.0$ $\mathbf{26.3}$ $\mathbf{26.3}$
-4 High contrast dot $19.6$ $7462.7$ $59.0$ $19.5$ $\mathbf{19.6}$ N/A $\mathbf{19.6}$
-5 Fidget spinner $4.7$ $451.2$ $5.0$ $19.1$ $1$.11111.11111.11111.1 N/A $\mathbf{4.7}$
-6 Whirling spider $3.2$ $35.1$ $13.0$ $4.2$ $298.6$ $5.3$ $\mathbf{3.2}$
-7 Flashing LED $2000.0$ $1996.0$ $1999.4$ $2000.1$ $\mathbf{2000.0}$ $1984.1$ $\mathbf{2000.0}$
-8 Refreshing mobile phone screen $240.0$ $2770.1$ $\mathbf{240.0}$ $\mathbf{240.0}$ $239.8$ $239.9$ $238.1$
-9 Speaker diaphragm $98.0$ $21.0$ $98.1$ $97.8$ $99.0$ $95.3$ $\mathbf{98.0}$
-10 Vibrating motor $40.0$ $52.2$ $\mathbf{40.0}$ $\mathbf{40.0}$ $370.4$ N/A $\mathbf{40.0}$
-11 Bike chain (side) $28.7$ $123.8$ $1.0$ $33.9$ $1$.11111.11111.11111.1 $44.7$ $\mathbf{28.7}$
-12 Bike chain (top) $23.0$ $51.6$ $22.0$ $30.8$ $1428.6$ $22.7$ $\mathbf{23.1}$
-Table 3: Frequency estimates (Hz) on the 12 sequence dataset.
-Methods: Simple baseline (Sec.
-3.1), FFT baseline (Sec.
-3.2), Prophesse Vibration Estimation [14], EB-ASM [2], Frequency-cam [13] and EE3P3D (Sec.
-3.3).
-Method Simple base. FFT base. Prophesee Vibr. Est. EB-ASM Freq-cam EE3P3D (ours)
-Felt disc with high contrast line $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$ $<\epsilon$
-Velcro disc (front) $14.6$ $0.5$ $<\epsilon$ $29520.9$ $0.5$ $<\epsilon$
-Velcro disc (side) $0.4$ $101.5$ $<\epsilon$ $9173.8$ $<\epsilon$ $<\epsilon$
-High contrast dot $37975.0$ $201.0$ $0.5$ $<\epsilon$ N/A $<\epsilon$
-Fidget spinner $9500.0$ $6.3$ $306.3$ $23540.4$ N/A $<\epsilon$
-Whirling spider $996.9$ $306.3$ $31.3$ $9231.3$ $65.6$ $<\epsilon$
-Flashing LED $0.2$ $0.03$ $0.005$ $<\epsilon$ $0.7$ $<\epsilon$
-Refreshing mobile phone screen $1054.2$ $<\epsilon$ $<\epsilon$ $0.08$ $0.04$ $0.8$
-Speaker diaphragm $78.6$ $0.1$ $0.2$ $1.0$ $2.8$ $<\epsilon$
-Vibrating motor $30.5$ $<\epsilon$ $<\epsilon$ $826.0$ N/A $<\epsilon$
-Bike chain (side) $331.4$ $96.5$ $18.1$ $3771.4$ $55.8$ $<\epsilon$
-Bike chain (top) $124.3$ $4.3$ $33.9$ $6111.3$ $1.3$ $0.4$
-Mean relative error $4175.5$ $59.7$ $32.5$ $6848.0$ $14.0$ $0.1$
-Max relative error $37975.0$ $306.2$ $306.3$ $29520.9$ $65.6$ $0.8$
-Table 4: Relative errors (%) on the 12 sequence dataset.
-Methods: Simple baseline (Sec.
-3.1), FFT baseline (Sec.
-3.2), Prophesse Vibration Estimation [14], EB-ASM [2], Frequency-cam [13] and EE3P3D (Sec.
-3.3), $\epsilon=10^{-3}$.
 6 Conclusion
-In this work, a novel contactless method for measuring the frequency of periodic phenomena is presented.
-The method makes only a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream.
-The proposed method, EE3P3D, is evaluated on twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), with targets lacking prominent features, being behind transparent materials, and having centrosymmetric shapes, achieving the mean relative error of $0.1\%$.
+A novel contactless method for measuring the rate of periodic phenomena has been presented.
+The method makes a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream.
+The method, EEPPR, was evaluated on a new dataset of twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), with targets lacking prominent features, being behind transparent materials, and having centrosymmetric shapes, achieving the mean relative error of $0.1\%$.
 The method works in non-frontal camera placement, measuring frequencies as high as $2000$ Hz (equivalent to 120 000120000120\,000120 000 RPM) even when a significant amount of noise is present.
-The dataset and codes are publicly available on GitHub.
+Limitations
+The method assumes the location where the period phenomenon is occurring is static. Handling of tracking and the template update is future work.
 Appendix 0.A Experimental Setup
-For the ’Felt disc with a high-contrast line’ experiment (see Sec. 4.2.1), we captured a power drill spinning a white felt disc marked with a black line, see Fig. 0.A.5.
+For the ’Felt disc with a high-contrast line’ experiment (see Sec. 5.2), we captured a power drill spinning a white felt disc marked with a black line, see Fig. 0.A.5.
 [Uncaptioned image] [Uncaptioned image]
@@
 Figure 0.A.5: Felt disc with a high-contrast line (see Sec.
-4.2.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($655\times 655$px) shown in (d).
+5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($655\times 655$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
 For the ’Fronto-parallel velcro disc’ experiment (see Sec.
-4.2.2), a drill spun the disc at $1266$ RPM with the event camera facing it directly, while a laser tachometer measured speed from the side of the disc where a reflective sticker was attached (Fig.
+5.2), a drill spun the disc at $1266$ RPM with the event camera facing it directly, while a laser tachometer measured speed from the side of the disc where a reflective sticker was attached (Fig.
 0.A.6a).
@@
 Figure 0.A.6: Fronto-parallel velcro disc (see Sec.
-4.2.2): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($713\times 713$px) shown in (d).
+5.2): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($713\times 713$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
-In the next experiment (’Velcro disc (side)’, see Sec. 4.2.3), the velcro disc was captured by the event camera through a glass sheet at $45\degree$ camera angle (see Fig. 4a).
+In the next experiment (’Velcro disc (side)’, see Sec. 5.2), the velcro disc was captured by the event camera through a glass sheet at $45\degree$ camera angle (see Fig. 4a).
 For the following experiment (’High contrast dot’), a sequence from Prophesee’s dataset was used, capturing an orbiting dark dot (see Sec.
-4.2.4).
+5.2).
 The reference photo was generated using Prophesee’s Event to Video method (see Fig.
@@
 Figure 0.A.7: High-contrast dot (see Sec.
-4.2.4): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($263\times 263$px) shown in (d).
+5.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($263\times 263$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
-This experiment (see Sec. 4.2.5) used event data from Prophesee’s dataset, capturing a rotating three-blade hand fidget spinner on a flat surface, see Fig. 0.A.8.
+This experiment (see Sec. 5.2) used event data from Prophesee’s dataset, capturing a rotating three-blade hand fidget spinner on a flat surface, see Fig. 0.A.8.
 [Uncaptioned image] [Uncaptioned image]
@@
 Figure 0.A.8: Hand fidget spinner (see Sec.
-4.2.5): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($239\times 239$px) shown in (d).
+5.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($239\times 239$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
 The sequence used for the next experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider (see Sec.
-4.2.6).
+5.2).
 A second rotating pattern is visible, caused by the shadow of the spider (see Fig.
@@
 Figure 0.A.9: Whirling Pholcus phalangioides (see Sec.
-4.2.6): (a) physical setup; (b) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($221\times 221$px) shown in (c); (e) events from a 250-millisecond window visualised in spatio-temporal space.
+5.2): (a) physical setup; (b) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($221\times 221$px) shown in (c); (e) events from a 250-millisecond window visualised in spatio-temporal space.
 Positive events are represented by white colour, and negative events are bright blue.
-For the LED experiment (see Sec. 4.3.1), a circuit with a light-emitting diode (see Fig. 0.A.10a) was used.
+For the LED experiment (see Sec. 5.3), a circuit with a light-emitting diode (see Fig. 0.A.10a) was used.
 [Uncaptioned image] [Uncaptioned image]
@@
 Figure 0.A.10: Flashing LED (see Sec.
-4.3.1): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($126\times 126$px) shown in (d).
+5.3): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($126\times 126$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
 The sequence used for the ’Refreshing mobile phone screen’ experiment (see Sec.
-4.3.2) captured the flicker of a mobile phone screen displaying a white colour at maximum brightness.
+5.3) captured the flicker of a mobile phone screen displaying a white colour at maximum brightness.
 A moving top-to-bottom pattern was visible when inspecting the event data (see Fig.
@@
 Figure 0.A.11: Refreshing mobile phone screen (see Sec.
-4.3.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within hardware-level region of interest and a time window equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($54\times 54$px) shown in (d).
+5.3): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within hardware-level region of interest and a time window equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($54\times 54$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
-In the following experiment (see Sec. 4.4.1), we captured a vibrating diaphragm of a speaker with two large low-frequency drivers, see Fig. 0.A.12a.
+In the following experiment (see Sec. 5.4), we captured a vibrating diaphragm of a speaker with two large low-frequency drivers, see Fig. 0.A.12a.
 [Uncaptioned image] [Uncaptioned image]
@@
 Figure 0.A.12: Speaker diaphragm (see Sec.
-4.4.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($631\times 631$px) shown in (d).
+5.4): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($631\times 631$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
-For the ’Vibrating motor’ experiment (see Sec. 4.4.2) we used a sequence from Prophesee’s dataset capturing a motor vibrating vertically at $40$ Hz; see Fig. 0.A.13.
+For the ’Vibrating motor’ experiment (see Sec. 5.4) we used a sequence from Prophesee’s dataset capturing a motor vibrating vertically at $40$ Hz; see Fig. 0.A.13.
 [Uncaptioned image] [Uncaptioned image]
@@
 Figure 0.A.13: Vibrating motor (see Sec.
-4.4.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($148\times 148$px) shown in (d).
+5.4): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($148\times 148$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
-The sequence used for this experiment (see Sec. 4.5.1) captured a bicycle chain from a close-up side view (see Fig. 0.A.14a).
+The sequence used for this experiment (see Sec. 5.5) captured a bicycle chain from a close-up side view (see Fig. 0.A.14a).
 [Uncaptioned image] [Uncaptioned image]
@@
 Figure 0.A.14: Bike chain from side view (see Sec.
-4.5.1): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within a hardware-level region of interest and a time window equal to the period of the observed phenomenon, highlighted region of interest ($204\times 204$px) shown in (d).
+5.5): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within a hardware-level region of interest and a time window equal to the period of the observed phenomenon, highlighted region of interest ($204\times 204$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.
-The sequence for the final experiment (see Sec. 4.5.2) captured the same bike chain from a top view perspective (see Fig. 0.A.15a).
+The sequence for the final experiment (see Sec. 5.5) captured the same bike chain from a top view perspective (see Fig. 0.A.15a).
 [Uncaptioned image] [Uncaptioned image]
@@
 Figure 0.A.15: Bike chain from a top view (see Sec.
-4.5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of lenght equal to the period of the observed phenomenon, highlighted region of interest ($180\times 180$px) shown in (d).
+5.5): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of lenght equal to the period of the observed phenomenon, highlighted region of interest ($180\times 180$px) shown in (d).
 Positive events are represented by white colour, and negative events are bright blue.

### v3

@@
 Abstract
-We present a novel method for measuring the period of phenomena like rotation, flicker and vibration, by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution.
+We present a novel method for measuring the rate of periodic phenomena (e.g., rotation, flicker, and vibration), by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution.
 The approach assumes that for a periodic phenomenon, a highly similar set of events is generated within a spatio-temporal window at a time difference corresponding to its period.
@@
 flashing light and vibration, and periodic motion, e.g., rotation, ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM).
-EEPPR significantly outperforms published methods on this dataset, achieving the mean relative error of $0.1\%$.
-The dataset and codes are available on GitHub.
+EEPPR significantly outperforms published methods on this dataset, achieving a mean relative error of $0.1\%$ setting new state of the art.
+The dataset and codes are publicly available on GitHub.
 keywords:
@@
 Precise quantification of rotational speed, for instance, finds applications in (i) monitoring rotating machinery components for performance evaluation and quality control [1], (ii) ensuring flight stability and manoeuvrability in drones [2], (iii) analysing sports equipment like ball tracking [3, 4], and (iv) optimising energy production in wind turbines [5].
-Traditional methods for measuring periodic phenomena often involve contact-based devices such as mechanical tachometers and rotary encoders.
+Traditional methods for measuring periodic phenomena often involve contact-based devices such as rotary encoders and mechanical tachometers.
 These approaches, while established, have inherent limitations: (i) physical contact with the target object can influence its movement and introduce measurement inaccuracies, (ii) contact methods may not be feasible in scenarios with delicate objects, confined spaces, or situations where interfering with the target significantly affects its movement.
@@
 Although contactless event-based methods exist, they are either handcrafted for a specific type of periodic phenomena [6] or require user supervision [7, 8].
-Figure 1: EEPPR method: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a 3D correlation of the template with the event stream is computed, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows.
 In this paper, we propose a non-contact method for measuring the properties of any periodic phenomena using an event camera.
-The input of our method is the event stream, and the output is a single scalar – the rate of the periodic phenomena.
-The method operates remotely, without requiring any modification of the observed object (e.g.
-a marker).
+The input of our method is the event stream, and the output is a single scalar, the rate of the periodic phenomena calculated as a robust point estimate within a spatial window.
+The method operates remotely, without requiring any modification of the observed object (e.g., a marker).
 The approach assumes that similar sets of pixel activations of an event camera (events) will occur at regular time intervals in any sequence capturing a periodic phenomenon.
 These intervals correspond to the period.
+For phenomena where the time of repetition varies, the method provides the time of each period which can be further analysed.
 The method was evaluated on a new dataset of twelve sequences of four types of periodic phenomena: (i) flickering light, (ii) object vibration, (iii) rotation, and (iv) repetitive translational movement.
@@
 3).
-The point estimate for the period lengths is obtained by calculating a median in a spatial window for robustness.
-For phenomena where the time of repetition varies, the method provides the time of each period which can be further analysed.
+(a) EEPPR method overview
+(b) Normalised correlation responses
+(c) Ground-truth rates
+Figure 1: (a) EEPPR: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a correlation of the template with the event stream is computed in 3D, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows.
+(b) Normalised correlation responses of a selected 3D template with $1000$ ms of spatio-temporal event stream.
+Periodic peaks are highly distinctive, highly regular and indicate a periodic phenomenon.
+(c) Ground-truth rates of experiments in order of appearance in this work.
 2 Related work
@@
 The time deltas between zero-level crossing are then used to estimate the rotation speed.
+In [15], two vibration estimation methods using high-contrast markers are proposed.
+Both apply the Fast Fourier Transform (FFT) to identify the dominant frequency.
+The first analyses marker displacement, achieving a maximum relative error of $1.43\%$ for vibrations up to $110$ Hz, the second analyses event generation rate variation within a bounding box, which moves with the marker’s centroid, with a maximum relative error of $1\%$ for vibrations up to $190$ Hz.
 In [4], a method for estimation of a table tennis ball spin was proposed.
@@
 Although their approach is unknown, we present the results of their method in our experiments.
+Table 1: EEPPR, average and maximum relative errors (%) as functions of window size on the dataset 5.1.
 Window size (px) $30\times 30$ $45\times 45$ $60\times 60$ $75\times 75$
-Avg relative error $37.78$ $\mathbf{19.34}$ $42.96$ $487.98$
-Max relative error $\mathbf{3225.00}$ $3362.50$ $7712.50$ $22737.50$
-Table 1: EEPPR, average and maximum relative errors (%) as functions of window size on the dataset 5.1.
+[…]
+[…]
+Table 2: EEPPR, dependence on template event count (see Sec. 4) on the dataset 5.1.
 Template ev count $100-1400$ $1500-3000$
@@
 Max relative error $22737.50$ $\mathbf{269.36}$
-Table 2: EEPPR, dependence on template event count (see Sec. 4) on the dataset 5.1.
 3 Baselines
@@
   1. 1.
-    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform with the Fast Fourier Transform (FFT)  [3] with the event creation time represented by a Dirac pulse.
+    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform [17] with the event creation time represented by a Dirac pulse.
   2. 2.
@@
 4 EEPPR – The proposed method
-Our method (see Fig. 1) detects periodic similarities along the time axis of event data in four steps:
+Our method (see Fig. 1(a)) detects periodic similarities along the time axis of event data in four steps:
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+(a) Physical setup           (b) XYTime visualisation (c) Target close-up          (d) Aggregated events
+Figure 2: Velcro disc with a non-frontal camera behind a glass sheet (see Sec.
+5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon.
+Positive events are represented by white colour, negative events are bright blue.
   1. 1.
@@
   2. 2.
-    Correlation template selection – the array $\mathbf{E}$ is split into a grid of $K$ spatially non-overlapping areas $\mathbf{A}_{i}$, $i\in\{1,\dots K\}$, of size $W\times W$.
+    Correlation template selection – the array $\mathbf{E}$ is split into a grid of $K$ spatially non-overlapping areas $\mathbf{A}_{i}$, $i~{}\in~{}\{1,\dots K\}$, of size $W\times W$.
 In each area, the depth (time axis size) of the correlation template of the area $\mathbf{T}_{i}$ is found by thresholding on a predefined number of events $N$.
@@
     Correlation in 3D – a correlation in 3D of the template $\mathbf{T}_{i}$ is computed with the area $\mathbf{A}_{i}$ (see Fig.
-2).
+1(b)).
 The time difference between consecutive peaks in correlation responses is the estimated period of the phenomenon.
@@
     Results aggregation – the output of the method is the median of estimated periods from all areas $\mathbf{A}_{i}$.
-Figure 2: Normalised correlation responses of a selected 3D template with $1000$ ms of spatio-temporal event stream.
-Periodic peaks are highly distinctive, highly regular and indicate a periodic phenomenon.
 5 Experiments
@@
 Using 3-fold cross-validation on the introduced dataset, the best accuracy is achieved using the window size $W$ set to $45\times 45$ px and the template event count $N$ set to $1800$ events.
-These parameters are proposed as default parameters of the method.
+We propose these parameter as default.
 We compare the results of the proposed method both with the baseline methods (see Sec.
@@
 We averaged all tachometer readings captured in one second ($3$ to $5$) to obtain a single reference value for comparison with the results of all methods.
-Figure 3: Ground-truth rates of experiments in order of appearance in this work.
 However, acquiring GT data this way was not feasible for all experiments.
@@
 The final dataset contains sequences of period phenomena with rates ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM), see Fig.
-3.
+1(c).
 Due to the non-constant rates of some periodic phenomena, we restricted our analysis to one-second segments of the event stream.
@@
 The Prophesee’s and our proposed method achieved the highest accuracy (relative error $<0.001$%), demonstrating robustness in this low-feature scenario.
-[Uncaptioned image] [Uncaptioned image]
-(a) Physical setup (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Target close-up (d) Aggregated events
-Figure 4: Velcro disc with a non-frontal camera behind a glass sheet (see Sec.
-5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon.
-Positive events are represented by white colour, negative events are bright blue.
 Velcro disc with a non-frontal camera behind a glass sheet
@@
 The velcro disc was spun at $1578$ RPM and captured at a $45\degree$ camera angle by the event camera through a glass sheet (see Fig.
-4), simulating a possible industry application when the observed object is a rotating machine part and a transparent visor protects the camera and machine operators.
+2), simulating a possible industry application when the observed object is a rotating machine part and a transparent visor protects the camera and machine operators.
 Interestingly, the Simple baseline method estimated the rate very well ($0.38\%$ relative error).
@@
 Refreshing screen
-The sequence used for this experiment captured the flicker of a mobile phone screen that displayed a white colour at maximum brightness.
-A top-to-bottom pattern was visible when inspecting the event data, likely due to Pulse Width Modulation (PWM) for backlight control.
+For this experiment we captured the flicker of a mobile phone screen displaying a white colour at maximum brightness.
+A top-to-bottom pattern is visible in the event data, likely due to Pulse Width Modulation (PWM) for backlight control.
 The GT rate was estimated manually, using the EE3P method and confirmed by findings in a study on the display of this phone [^7] Available at https://bit.ly/notebookcheck-SamsungGalaxyS21UltraReview..
@@
 3).
+Table 3: Frequency estimates (Hz) and relative errors (%) on the 12 sequence dataset.
+Methods: Simple baseline (Sec.
+3), FFT baseline (Sec.
+3), Prophesee Vibration Estimation [8], EB-ASM [6], Frequency-cam [7] and EEPPR (Sec.
+4).
 Frequency Estimates (Hz) Relative Errors (%)
@@
 12 Bike chain (top) 23.0 51.6 22.0 30.8 1428.6 22.7 23.1 124.3 4.3 33.9 6111.3 1.3 0.4
-Table 3: Frequency estimates (Hz) and relative errors (%) on the 12 sequence dataset.
-Methods: Simple baseline (Sec.
-3), FFT baseline (Sec.
-3), Prophesee Vibration Estimation [14], EB-ASM [1], Frequency-cam [13] and EEPPR (Sec.
-4).
+Mean relative error $4175.5$ $59.7$ $32.5$ $6848.0$ $14.0$ $0.1$
+Max relative error $37975.0$ $306.3$ $306.3$ $29520.9$ $65.6$ $0.8$
 5.4 Measuring vibration frequency of a speaker and a motor
 Speaker diaphragm
-In this experiment, we captured and analysed a sequence capturing a vibrating diaphragm of a speaker with two large low-frequency drivers.
+In this experiment, we captured and analysed vibrating diaphragm of a speaker with two large low-frequency drivers.
 An Android application [^8] Application available at https://bit.ly/Frequency-Sound-Generator-APK.
@@
 The proposed method was the only method that achieved a relative error lower than $0.1\%$.
-The FFT baseline and Prophesee methods achieved relative error under $0.2\%$, followed by EB-ASM ($1.02\%$) and Frequency-cam ($2.8\%$).
-The Simple baseline method deviated the most from the reference frequency ($78.571\%$ relative error); see Tab.
+The FFT baseline and Prophesee methods achieved relative error under $0.2\%$, followed by EB-ASM ($1.0\%$) and Frequency-cam ($2.8\%$).
+The Simple baseline method deviated the most from the reference frequency ($78.6\%$ relative error); see Tab.
 3.
@@
 The experiments where the target lacked high-contrast features were shown to be the most challenging.
-The worst performing methods were the Simple baseline method and the EB-ASM method [1] achieving MRE of $4175.511\%$ and $6848.011\%$ respectively (see Tab.
+The worst performing methods were the Simple baseline method and the EB-ASM method [6] achieving MRE of $4175.5\%$ and $6848.0\%$ respectively (see Tab.
 3).
-The third lowest mean relative error across all experiments ($32.532\%$) was achieved by the Vibration Estimation method by Prophesee.
-The Frequency-cam method achieved the second lowest mean relative error (MRE) of $14.083\%$, which is significantly higher than the MRE achieved by the proposed method ($0.1\%$).
+The third lowest mean relative error across all experiments ($32.5\%$) was achieved by the Vibration Estimation method by Prophesee.
+The Frequency-cam method achieved the second lowest mean relative error (MRE) of $14.1\%$, which is significantly higher than the MRE achieved by the proposed method ($0.1\%$).
 Note that the Frequency-cam method could not analyse three of the twelve sequences we used in this work, which originated from the public dataset by Prophesee due to an older file encoding type.
@@
 6 Conclusion
-A novel contactless method for measuring the rate of periodic phenomena has been presented.
-The method makes a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream.
-The method, EEPPR, was evaluated on a new dataset of twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), with targets lacking prominent features, being behind transparent materials, and having centrosymmetric shapes, achieving the mean relative error of $0.1\%$.
+A novel contactless event-based method, EEPPR, for measuring the rate of various periodic phenomena, such as rotational speed, vibration, flicker and periodic movement, was presented.
+The method makes only a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream.
+The rate of periodic phenomena is computed by measuring time deltas between peaks detected in correlation responses of an automatically selected template correlated with a spatio-temporal event stream.
+EEPPR was evaluated on a new proposed dataset of twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), including targets lacking prominent features, being behind transparent materials, or having centrosymmetric shapes, achieving the mean relative error of $0.1\%$ setting new state of the art.
 The method works in non-frontal camera placement, measuring frequencies as high as $2000$ Hz (equivalent to 120 000120000120\,000120 000 RPM) even when a significant amount of noise is present.
+The dataset and code are publicly available on GitHub.
 Limitations
-The method assumes the location where the period phenomenon is occurring is static. Handling of tracking and the template update is future work.
-Appendix 0.A Experimental Setup
-For the ’Felt disc with a high-contrast line’ experiment (see Sec. 5.2), we captured a power drill spinning a white felt disc marked with a black line, see Fig. 0.A.5.
-[Uncaptioned image] [Uncaptioned image]
-(a) Physical setup (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.5: Felt disc with a high-contrast line (see Sec.
-5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($655\times 655$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-For the ’Fronto-parallel velcro disc’ experiment (see Sec.
-5.2), a drill spun the disc at $1266$ RPM with the event camera facing it directly, while a laser tachometer measured speed from the side of the disc where a reflective sticker was attached (Fig.
-0.A.6a).
-[Uncaptioned image] [Uncaptioned image]
-(a) Physical setup (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) RoI
-Figure 0.A.6: Fronto-parallel velcro disc (see Sec.
-5.2): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($713\times 713$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-In the next experiment (’Velcro disc (side)’, see Sec. 5.2), the velcro disc was captured by the event camera through a glass sheet at $45\degree$ camera angle (see Fig. 4a).
-For the following experiment (’High contrast dot’), a sequence from Prophesee’s dataset was used, capturing an orbiting dark dot (see Sec.
-5.2).
-The reference photo was generated using Prophesee’s Event to Video method (see Fig.
-0.A.7a).
-[Uncaptioned image] [Uncaptioned image]
-(a) Target visualisation (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.7: High-contrast dot (see Sec.
-5.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($263\times 263$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-This experiment (see Sec. 5.2) used event data from Prophesee’s dataset, capturing a rotating three-blade hand fidget spinner on a flat surface, see Fig. 0.A.8.
-[Uncaptioned image] [Uncaptioned image]
-(a) Target visualisation (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.8: Hand fidget spinner (see Sec.
-5.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($239\times 239$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-The sequence used for the next experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider (see Sec.
-5.2).
-A second rotating pattern is visible, caused by the shadow of the spider (see Fig.
-0.A.9d,e).
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-(a) Physical setup                 (b) Target close-up (c) Region of Interest
-[Uncaptioned image] [Uncaptioned image]
-(d) Aggregated events (e) XYTime visualisation
-Figure 0.A.9: Whirling Pholcus phalangioides (see Sec.
-5.2): (a) physical setup; (b) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($221\times 221$px) shown in (c); (e) events from a 250-millisecond window visualised in spatio-temporal space.
-Positive events are represented by white colour, and negative events are bright blue.
-For the LED experiment (see Sec. 5.3), a circuit with a light-emitting diode (see Fig. 0.A.10a) was used.
-[Uncaptioned image] [Uncaptioned image]
-(a) Physical setup (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.10: Flashing LED (see Sec.
-5.3): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($126\times 126$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-The sequence used for the ’Refreshing mobile phone screen’ experiment (see Sec.
-5.3) captured the flicker of a mobile phone screen displaying a white colour at maximum brightness.
-A moving top-to-bottom pattern was visible when inspecting the event data (see Fig.
-0.A.11b).
-[Uncaptioned image] [Uncaptioned image]
-(a) Physical setup (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.11: Refreshing mobile phone screen (see Sec.
-5.3): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within hardware-level region of interest and a time window equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($54\times 54$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-In the following experiment (see Sec. 5.4), we captured a vibrating diaphragm of a speaker with two large low-frequency drivers, see Fig. 0.A.12a.
-[Uncaptioned image] [Uncaptioned image]
-(a) Physical setup (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.12: Speaker diaphragm (see Sec.
-5.4): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($631\times 631$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-For the ’Vibrating motor’ experiment (see Sec. 5.4) we used a sequence from Prophesee’s dataset capturing a motor vibrating vertically at $40$ Hz; see Fig. 0.A.13.
-[Uncaptioned image] [Uncaptioned image]
-(a) Target visualisation (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.13: Vibrating motor (see Sec.
-5.4): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($148\times 148$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-The sequence used for this experiment (see Sec. 5.5) captured a bicycle chain from a close-up side view (see Fig. 0.A.14a).
-[Uncaptioned image] [Uncaptioned image]
-(a) Physical setup (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.14: Bike chain from side view (see Sec.
-5.5): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within a hardware-level region of interest and a time window equal to the period of the observed phenomenon, highlighted region of interest ($204\times 204$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
-The sequence for the final experiment (see Sec. 5.5) captured the same bike chain from a top view perspective (see Fig. 0.A.15a).
-[Uncaptioned image] [Uncaptioned image]
-(a) Physical setup (b) XYTime visualisation
-[Uncaptioned image] [Uncaptioned image]
-(c) Aggregated events (d) Region of Interest
-Figure 0.A.15: Bike chain from a top view (see Sec.
-5.5): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of lenght equal to the period of the observed phenomenon, highlighted region of interest ($180\times 180$px) shown in (d).
-Positive events are represented by white colour, and negative events are bright blue.
+The method assumes the periodic phenomenon undergoes only periodic movement without additional motion.
+Handling of tracking and the template update is future work.
+Aggregating events within fixed intervals $t_{\text{quant}}$ restricts both the maximal measurable frequency and the theoretical limits of precision.
+Reducing the aggregation length $t_{\text{quant}}$ mitigates this at a cost of higher computation time.
+Future work will explore sparse data correlation to avoid aggregation.
+Acknowledgements
+The research reported in this paper has been funded by BMK, BMAW, and the State of Upper Austria within the SCCH competence center INTEGRATE (grant no.
+892418) part of the FFG COMET Competence Excellent Technologies Programme, and by the Czech Technical University in Prague grant No.
+SGS23/173/OHK3/3T/13.